# TabTreeFormer: Tabular Data Generation Using Hybrid Tree-Transformer

### v2

@@ 3.3 Auto-regressive Transformer
TabTreeFormer is simply trained as an auto-regressive transformer. The input token sequence is constructed by concatenating the outcome from the tree model and tokenizer. Totally 5 major types of tokens are required and the vocabulary size is $V=n_{l}+n_{c}+n_{b}+n_{q}+3$. The number of tokens in a constructed sequence of tokens corresponding to one row is $L=2+T+m_{d}+2m_{c}$, including BOS and EOS tokens. Both of them are much lower than the need for natural language models. A shared set of leaf tokens across all trees and similarly shared sets of categorical, bin, and quantile tokens across all features are used. Distinctions between different trees and features are effectively encoded through positional information, leveraging the positional embeddings inherent in most transformers.
A key observation on the construction of token sequence is that the valid range of tokens at each position is fixed for a given table. Consequently, tokens can be sampled exclusively from this precomputed set of valid options during generation. This ensures that every generated token sequence corresponds to a valid row, eliminating the need for rejecting invalid samples – a common practice in existing methods (Borisov et al., 2023; Solatorio & Dupriez, 2023; Zhao et al., 2023) – and significantly accelerating the sampling process.
Additionally, transformers are prone to memorization, particularly when trained on tabular data, where datasets are much smaller compared to typical natural language corpora (Borisov et al., 2023). To avoid over-fitting in terms of memorization, we heavily mask the input and maintain the target output unmasked. The rest of the training and generation settings are similar to the training and generation of classical causal language models, except for a modified loss function to account for the ordinal relations between quantile tokens, with details introduced in Section 3.4.
@@ 3.4 Loss Function
Formally, let the predicted logits for a token be $\mathbf{z}\in\mathbb{R}^{V}$, and let $t\in\{1,\dots,V\}$ be the target label. Recall the classical definition of CEL as given in Equation 1.
$\mathcal{L}_{\text{ce}}(\mathbf{z},t)=-\log\frac{e^{z_{t}}}{\sum_{i=1}^{V}e^{z_{i}}}$ (1)
@@
If $w_{ti}\geq 0$ is a function of the distance of the current class $i$ to the target class $t$, namely, $w_{ti}=f(|t-i|)$ where $f:\mathbb{N}\rightarrow\mathbb{R}_{+},|t-i|\mapsto w_{i}$, and $f$ is monotonically increasing, then the OCEL penalizes adjacent classes less than farther classes.
Both Lemma 1 and Theorem 1 are intuitive, and we provide rigorous proofs in Appendix A.1-A.2.
Theorem 1 requires the definition of $w_{ti}=f(|t-i|)$ that satisfies the monotonically increasing constraint to be used for OCEL. Any satisfying definition works technically. In this paper, we define the weight $w_{ti}$ for $z_{i}$ to be obtained by Equation 3.
@@
Note that not all tokens are part of the ordinal relation, so the OCEL is applied only to valid quantile tokens at the corresponding positions. To enable the model to learn to generate valid quantile tokens, we introduce an additional modified CEL to distinguish valid tokens from invalid ones, which complements OCEL. Formally, let valid quantile tokens have IDs in the range $[Q_{s},Q_{e})$ The valid token group loss (for quantile tokens at this position) is defined by Equation 4 when $t\in[Q_{s},Q_{e})$.
@@ 4.1 Experimental Setup
Baseline models. We compare the performance of TabTreeFormer against several influential and high-performing methods in tabular data generation including non-neural networks, GANs, VAEs, diffusion models, and auto-regressive transformers: ARF (Watson et al., 2023), CTAB-GAN+ (CTAB+) (Zhao et al., 2024), CTGAN (Xu et al., 2019), TVAE (Xu et al., 2019), TabDDPM (Kotelnikov et al., 2023), TabSyn (Zhang et al., 2024), and GReaT (Borisov et al., 2023), REaLTabFormer (RTF) (Solatorio & Dupriez, 2023). Implementation of baselines is described in Appendix C.3.
@@
  • Utility exhibits the quality of synthetic data by testing its performance on downstream tasks. We evaluate based on the established Train-on-Synthetic, Test-on-Real (TSTR) machine learning efficacy (MLE) evaluation framework (Xu et al., 2019), namely, synthetic data generated based on real training data is tested on a hold-out real test set. Three models are used: linear (i.e. linear and logistic regression; LN), random forest (RF) (Breiman, 2001), and XGBoost (XGB) (Chen & Guestrin, 2016). Better generative models tend to yield better scores on downstream tasks.
  • Privacy is a crucial criterion when the synthetic data serves the use case of privacy-preserving data sharing. We use the distance to the closest record (DCR) (Zhao et al., 2021) to evaluate the privacy of synthetic data. We compare the DCR from synthetically generated data and from hold-out real (test) data to the real training data and privacy preservation is demonstrated by DCRs calculated on the latter being no smaller than the former, tested by Mann-Whitney U Test (Mann & Whitney, 1947). Privacy is at risk of disclosure when $p$-values are smaller than 0.05.
@@ 4.2 Fidelity
Figure 4: Visualization of the marginal densities of the generated data of baseline tabular auto-regressive transformers and TabTreeFormer in comparison to the real data on some representative multimodal-distributed continuous columns. All these models use Distill-GPT2 (or its structure and hyperparameters) as backbone.
Figure 5: Performance comparison of TabTreeFormer versus top-4 performing baselines in fidelity on pair-wise correlation in 3 experimented datasets. Real (absolute) correlation values are presented on the left, and the absolute error in correlation values in synthetic data from different models, capped at 0.4 for better visibility, are shown at right (the darker, the better). The left-top features are categorical fatures, and right-bottom features are numeric features.
Figure 6: Real vs. synthetic abs. corr.
Figure 7: Feature count vs. “Trend” score improvement of TabTreeFormer from some baselines.
Marginal density distribution of columns (Shape). TabTreeFormer achieves comparable performance in “Shape” metric, as shown in Fig. 3. Meanwhile, it can be seen from Fig. 4 that TabTreeFormer captures better multimodal distribution compared to other auto-regressive transformers, which validates the effectiveness of our dual-quantization tokenizer design. More fidelity results can be found in Table 9 and Fig. 11 in Appendix D.1.
Pair-wise correlation between columns (Trend). TabTreeFormer outperforms all baselines in “Trend” metric with a significant margin, as presented in Fig. 3, implying the superior capability of TabTreeFormer in learning inter-feature relations. It can be observed from Fig. 7-7 that even strong baselines struggle with correlations involving categorical features and correlation values are not close to 0. Moreover, TabTreeFormer tends to show a more significant improvement from auto-regressive transformer baselines on datasets with a larger number of features as referred in Fig. 7. This also demonstrates the advantage of introducing the inductive bias on low-correlations so that the models learn relations among a larger number of features better by filtering out less correlated features.
@@ 4.3 Utility (MLE)
Table 2: MLE performance between different models. The top two rows calculate the global average scores and relative error (w.r.t real data). The middle three rows show the average performance in each ML model, and the bottom two rows show the average performances in classification (clf.) and regression (reg.). Classification tasks are evaluated by weighted AUC ROC scores, and regression tasks are evaluated by $R^{2}$ scores. Best scores and second best scores are highlighted in bold with and without underscore, respectively.
ARF CTAB+ CTGAN DDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
Avg. $0.843_{\pm 0.155}$ $0.680_{\pm 0.271}$ $0.810_{\pm 0.189}$ $0.552_{\pm 0.865}$ $0.800_{\pm 0.206}$ $0.866_{\pm 0.130}$ $0.865_{\pm 0.123}$ $0.813_{\pm 0.189}$ $0.863_{\pm 0.127}$ $\boldsymbol{0.878}_{\pm 0.113}$ $\underline{\boldsymbol{0.880}}_{\pm 0.111}$
RE $0.057_{\pm 0.077}$ $0.244_{\pm 0.274}$ $0.099_{\pm 0.137}$ $0.483_{\pm 1.214}$ $0.110_{\pm 0.170}$ $0.027_{\pm 0.040}$ $0.027_{\pm 0.029}$ $0.095_{\pm 0.142}$ $0.029_{\pm 0.040}$ $\boldsymbol{0.017}_{\pm 0.023}$ $\underline{\boldsymbol{0.016}}_{\pm 0.023}$
LN $0.835_{\pm 0.176}$ $0.686_{\pm 0.258}$ $0.810_{\pm 0.193}$ $0.570_{\pm 0.882}$ $0.800_{\pm 0.197}$ $0.851_{\pm 0.153}$ $0.849_{\pm 0.156}$ $0.814_{\pm 0.163}$ $0.853_{\pm 0.154}$ $\boldsymbol{0.861}_{\pm 0.145}$ $\underline{\boldsymbol{0.862}}_{\pm 0.143}$
RF $0.842_{\pm 0.156}$ $0.681_{\pm 0.273}$ $0.806_{\pm 0.200}$ $0.566_{\pm 0.910}$ $0.798_{\pm 0.220}$ $0.873_{\pm 0.122}$ $0.872_{\pm 0.109}$ $0.810_{\pm 0.210}$ $0.866_{\pm 0.121}$ $\boldsymbol{0.888}_{\pm 0.095}$ $\underline{\boldsymbol{0.891}}_{\pm 0.090}$
XGB $0.851_{\pm 0.147}$ $0.672_{\pm 0.307}$ $0.814_{\pm 0.194}$ $0.519_{\pm 0.897}$ $0.802_{\pm 0.222}$ $0.874_{\pm 0.125}$ $0.874_{\pm 0.110}$ $0.814_{\pm 0.209}$ $0.870_{\pm 0.116}$ $\boldsymbol{0.884}_{\pm 0.102}$ $\underline{\boldsymbol{0.886}}_{\pm 0.101}$
clf. $0.907_{\pm 0.074}$ $0.750_{\pm 0.227}$ $0.895_{\pm 0.065}$ $0.850_{\pm 0.114}$ $0.868_{\pm 0.130}$ $0.917_{\pm 0.064}$ $0.913_{\pm 0.071}$ $0.890_{\pm 0.081}$ $0.918_{\pm 0.062}$ $\underline{\boldsymbol{0.922}}_{\pm 0.063}$ $\boldsymbol{0.922}_{\pm 0.065}$
reg. $0.586_{\pm 0.122}$ $0.400_{\pm 0.266}$ $0.472_{\pm 0.127}$ $-0.641_{\pm 1.467}$ $0.528_{\pm 0.239}$ $0.660_{\pm 0.122}$ $0.674_{\pm 0.095}$ $0.506_{\pm 0.187}$ $0.643_{\pm 0.061}$ $\boldsymbol{0.698}_{\pm 0.083}$ $\underline{\boldsymbol{0.710}}_{\pm 0.092}$
The MLE results are presented in Table 2, whose raw scores are in Table 10 in Appendix D.2, show that TabTreeFormer outperforms baseline models steadily regardless of the complexity of the ML model used, task, and dataset. TabTreeFormer-S achieves similar performance as the best-performing baselines (REaLTabForme and TabSyn), and the M and L versions of TabTreeFormer outperforms them by a significant margin of around 40%, assessed by relative errors of ML performance compared to real data. Furthermore, out of the 30 MLE scores, TabTreeFormer-M or -L outperforms all other baselines in 20 cases (2/3), and they both outperform all other baselines in 16 cases ($\geq$1/2).
As an important foundation of the superior capability of tree-based models in capturing potentially low-correlations inductive bias, rotational invariance of models is an aspect worthwhile analyzing. As visualized in Fig. 9, rotationally invariant generators (e.g. CTAB-GAN+, CTGAN, GReaT, TVAE) are generally worse than non-rotationally invariant counterparts (e.g. REaLTabFormer and our TabTreeFormer by positional embedding in a fixed order and TabSyn by VAE-encoding of data). Also, TabTreeFormer’s design outperforms rotationally invariant baseline generators more on better-performing non-rotationally invariant downstream models (RF and XGB) than on rotationally invariant downstream models (LN). This again emphasizes the importance of applying non-rotationally invariant models to introduce tabular inductive bias (for low correlations).
@@ 4.4 Privacy (DCR)
Table 3: Summarized performance comparison between different models in terms of privacy. The first row shows the number of violations, and the second row shows the minimum $p$ value for privacy leakage risk, the closer the value to 0, the higher the risk of privacy leakage. Risky values are highlighted in red.
@@
As summarized in Table 3 (raw scores in Table 11 in Appendix D.3), TabTreeFormer-M and L are safe in all the 10 datasets. TabTreeFormer-S has a privacy leakage in only one dataset, with the $p$-value very close to 0.05. In comparison, GReaT has severe privacy concerns in 5 out of 10 datasets. Therefore, while auto-regressive transformers have privacy leakage concerns, TabTreeFormer and REaLTabFormer are generally privacy-preserving, validating the effectiveness of using masking in preserving privacy for tabular auto-regressive transformers.
4.5 Efficiency
@@
Observing results reported above, the performance of TabTreeFormer-M is better than all baselines in most metrics, while it has only about 5M parameters and its most competitive counterparts have more than 20M (TabSyn) or even more than 80M parameters (GReaT and REaLTabFormer). Also, as shown in Table 4, there is no significant performance difference between TabTreeFormer-M and TabTreeFormer-L, although their sizes differ by a factor of 8. This demonstrates the capability of TabTreeFormer in maintaining high efficiency without significant compromise in effectiveness.
Table 5: Training time (s) comparison of TabTreeFormer compared to other auto-regressive transformer baselines on large datasets. Best values are highlighted in bold with underscore, and second best values are highlighted in bold without underscore.

### v5

@@ 3.3 Auto-regressive Transformer
TabTreeFormer is simply trained as an auto-regressive transformer. The input token sequence is constructed by concatenating the outcome from the tree model and tokenizer. It requires 5 types of tokens and a vocabulary size of $V=n_{l}+n_{c}+n_{b}+n_{q}+3$. The number of tokens in a sequence corresponding to one row is $L=2+T+m_{d}+2m_{c}$, including BOS and EOS tokens. Both the vocabulary size and sequence length are much smaller than the need for natural language models. A shared set of leaf tokens across all trees and similarly shared sets of category, bin, and quantile tokens across all features are used. Distinctions between different trees and features are effectively encoded through positional information, leveraging the positional embeddings inherent in most transformers.
A key observation on the construction of token sequence is that the valid range of tokens at each position is fixed for a given table. Consequently, tokens can be sampled exclusively from this precomputed set of valid options during generation. This ensures that every generated token sequence corresponds to a valid row, eliminating the need for rejecting invalid sequences (Borisov et al., 2023; Solatorio & Dupriez, 2023; Zhao et al., 2023), thus significantly accelerating the sampling process.
Additionally, transformers are prone to memorization, particularly when trained on tabular data, where datasets are much smaller compared to typical natural language corpora (Borisov et al., 2023). To avoid memorization, we heavily mask the input and maintain the target output unmasked. The rest of the training and generation settings are similar to the training and generation of classical causal language models, except for a modified loss function to account for the ordinal relations between quantile tokens, with details introduced in Section 3.4.
@@ 3.4 Loss Function
Formally, let the predicted logits for a token be $\mathbf{z}\in\mathbb{R}^{V}$, the probability after softmax be $\mathbf{p}\in[0,1]^{V}$ ($\|\mathbf{p}\|_{1}=1$), and let $t\in\{1,\dots,V\}$ be the target label. Recall the classical definition of CEL as given in Equation 1.
$\mathcal{L}_{\text{ce}}(\mathbf{z},t)=-\log\frac{e^{z_{t}}}{\sum_{i=1}^{V}e^{z_{i}}}=-\log p_{t}\vspace{-0.2em}$ (1)
@@
Let $f(|t-i|)$ be a variant of the distance between current class $i$ and target class $t$ when $f$ is non-negative and monotonically increasing. Let the weighted sum of this variant of distances to target incurred by some logit $\mathbf{z}$ be $D(\mathbf{z})=\sum_{i=1}^{V}p_{i}w_{i}$. If $w_{ti}=f(|t-i|)$ for some $f$, then for $\mathbf{z},\mathbf{z}^{*}$ with $\mathcal{L}_{\text{ce}}(\mathbf{z},t)=\mathcal{L}_{\text{ce}}(\mathbf{z}^{*},t)$ and $D(\mathbf{z})<D(\mathbf{z}^{*})$, we must have $\mathcal{L}_{\text{oce}}(\mathbf{z},t)<\mathcal{L}_{\text{oce}}(\mathbf{z}^{*},t)$.
Both Lemma 1 and Theorem 1 are intuitive, and we provide rigorous proofs in Appendix A.1-A.2. Theorem 1 implies that the OCEL penalizes (value being smaller) classes closer to the target class less than farther ones. Any definition of $w_{ti}=f(|t-i|)$ satisfying the monotonicity constraint works for OCEL technically. In this paper, we define the weight $w_{ti}$ for $z_{i}$ to be obtained by Equation 3.
@@
Note that not all tokens are part of the ordinal relation, so the OCEL is applied only to valid quantile tokens at the corresponding positions. To enable the model to learn to generate valid quantile tokens, we introduce an additional modified CEL to distinguish valid tokens from invalid ones, which complements OCEL. Formally, let valid quantile tokens have IDs in the range $[Q_{s},Q_{e})$. The valid token group loss (for quantile tokens at this position) is defined by Equation 4 when $t\in[Q_{s},Q_{e})$.
@@ 4 Experiments
Figure 4: Comparison of marginal densities for generated data from baseline tabular auto-regressive transformers and TabTreeFormer against real data on representative multimodal continuous columns. All models use Distill-GPT2 as the backbone.
@@ 4.1 Experimental Setup
Baseline models. We compare the performance of TabTreeFormer against several influential and high-performing methods in tabular data generation including non-neural networks, GANs, VAEs, diffusion models, and auto-regressive transformers: ARF (Watson et al., 2023), CTAB-GAN+ (CTAB+) (Zhao et al., 2024), CTGAN (Xu et al., 2019), TVAE (Xu et al., 2019), TabDDPM (Kotelnikov et al., 2023), TabSyn (Zhang et al., 2024), GReaT (Borisov et al., 2023) and REaLTabFormer (RTF) (Solatorio & Dupriez, 2023). Implementation of baselines is described in Appendix C.3.
Figure 5: Performance comparison of TabTreeFormer versus top-4 performing baselines in fidelity on pair-wise correlation in 3 experimented datasets. Real (absolute) correlation values are presented on the left, and the absolute error in correlation values in synthetic data from different models, capped at 0.4 for better visibility, are shown at the right (the darker, the better). The left-top features are categorical features, and the right-bottom features are numeric features.
Figure 6: Real vs. synthetic abs. corr.
Figure 7: Feature count vs. Trend score improvement of TabTreeFormer from baselines.
Table 2: MLE performance between different models. The top two rows calculate the global average scores and relative error (w.r.t real data). The middle three rows show the average performance in each ML model, and the bottom two rows show the average performances in classification (clf.) and regression (reg.). Classification tasks are evaluated by weighted AUC ROC scores, and regression tasks are evaluated by $R^{2}$ scores. The best scores and the second best scores are highlighted in bold with and without underscore, respectively.
ARF CTAB+ CTGAN TabDDPM GReaT RTF TabSyn TVAE TTF-S TTF-M TTF-L
Avg. ($\uparrow$) $0.843_{\pm 0.155}$ $0.680_{\pm 0.271}$ $0.810_{\pm 0.189}$ $0.552_{\pm 0.865}$ $0.800_{\pm 0.206}$ $0.866_{\pm 0.130}$ $0.865_{\pm 0.123}$ $0.813_{\pm 0.189}$ $0.863_{\pm 0.127}$ $\boldsymbol{0.878}_{\pm 0.113}$ $\underline{\boldsymbol{0.880}}_{\pm 0.111}$
RE ($\downarrow$) $0.057_{\pm 0.077}$ $0.244_{\pm 0.274}$ $0.099_{\pm 0.137}$ $0.483_{\pm 1.214}$ $0.110_{\pm 0.170}$ $0.027_{\pm 0.040}$ $0.027_{\pm 0.029}$ $0.095_{\pm 0.142}$ $0.029_{\pm 0.040}$ $\boldsymbol{0.017}_{\pm 0.023}$ $\underline{\boldsymbol{0.016}}_{\pm 0.023}$
LN ($\uparrow$) $0.835_{\pm 0.176}$ $0.686_{\pm 0.258}$ $0.810_{\pm 0.193}$ $0.570_{\pm 0.882}$ $0.800_{\pm 0.197}$ $0.851_{\pm 0.153}$ $0.849_{\pm 0.156}$ $0.814_{\pm 0.163}$ $0.853_{\pm 0.154}$ $\boldsymbol{0.861}_{\pm 0.145}$ $\underline{\boldsymbol{0.862}}_{\pm 0.143}$
RF ($\uparrow$) $0.842_{\pm 0.156}$ $0.681_{\pm 0.273}$ $0.806_{\pm 0.200}$ $0.566_{\pm 0.910}$ $0.798_{\pm 0.220}$ $0.873_{\pm 0.122}$ $0.872_{\pm 0.109}$ $0.810_{\pm 0.210}$ $0.866_{\pm 0.121}$ $\boldsymbol{0.888}_{\pm 0.095}$ $\underline{\boldsymbol{0.891}}_{\pm 0.090}$
XGB ($\uparrow$) $0.851_{\pm 0.147}$ $0.672_{\pm 0.307}$ $0.814_{\pm 0.194}$ $0.519_{\pm 0.897}$ $0.802_{\pm 0.222}$ $0.874_{\pm 0.125}$ $0.874_{\pm 0.110}$ $0.814_{\pm 0.209}$ $0.870_{\pm 0.116}$ $\boldsymbol{0.884}_{\pm 0.102}$ $\underline{\boldsymbol{0.886}}_{\pm 0.101}$
clf.($\uparrow$) $0.907_{\pm 0.074}$ $0.750_{\pm 0.227}$ $0.895_{\pm 0.065}$ $0.850_{\pm 0.114}$ $0.868_{\pm 0.130}$ $0.917_{\pm 0.064}$ $0.913_{\pm 0.071}$ $0.890_{\pm 0.081}$ $0.918_{\pm 0.062}$ $\underline{\boldsymbol{0.922}}_{\pm 0.063}$ $\boldsymbol{0.922}_{\pm 0.065}$
reg. ($\uparrow$) $0.586_{\pm 0.122}$ $0.400_{\pm 0.266}$ $0.472_{\pm 0.127}$ $-0.641_{\pm 1.467}$ $0.528_{\pm 0.239}$ $0.660_{\pm 0.122}$ $0.674_{\pm 0.095}$ $0.506_{\pm 0.187}$ $0.643_{\pm 0.061}$ $\boldsymbol{0.698}_{\pm 0.083}$ $\underline{\boldsymbol{0.710}}_{\pm 0.092}$
@@
  • Utility exhibits the quality of synthetic data by testing its performance on downstream tasks. We evaluate based on the established Train-on-Synthetic, Test-on-Real (TSTR) machine learning efficacy (MLE) evaluation framework (Xu et al., 2019), namely, synthetic data generated based on real training data is tested on a hold-out real test set. Three models are used: linear (i.e., linear and logistic regression; LN), random forest (RF) (Breiman, 2001), and XGBoost (XGB) (Chen & Guestrin, 2016). Classification performance is evaluated by weighted AUC ROC and regression is evaluated by $R^{2}$. For both metrics, better generative models tend to yield higher scores on downstream tasks.
  • Privacy is a crucial criterion when the synthetic data serves the use case of privacy-preserving data sharing. We use the distance to the closest record (DCR) (Zhao et al., 2021) to evaluate the privacy of synthetic data. We compare the DCR from synthetically generated data and from hold-out real (test) data to the real training data and privacy preservation is demonstrated by DCRs calculated on the latter being no smaller than the former, tested by Mann-Whitney U Test (Mann & Whitney, 1947). Privacy is at risk when $p$-values are smaller than 0.05.
@@ 4.2 Fidelity
Marginal density distribution of columns (Shape). TabTreeFormer achieves comparable performance to baseline models in “Shape” metric, as shown in Fig. 3. Meanwhile, it can be seen from Fig. 4 that TabTreeFormer captures better multimodal distribution compared to other auto-regressive transformers, which validates the effectiveness of our dual-quantization tokenizer design. More fidelity results can be found in Table 9 and Fig. 11 in Appendix D.1.
Pair-wise correlation between columns (Trend). TabTreeFormer outperforms all baselines in “Trend” metric with a significant margin, as presented in Fig. 3, implying the superior capability of TabTreeFormer in learning inter-feature relations. It can be observed from Fig. 7-7 that even strong baselines struggle with correlations involving categorical features and correlation values are not close to 0. Moreover, TabTreeFormer tends to show a more significant improvement from auto-regressive transformer baselines on datasets with a larger number of features as referred in Fig. 7. This also demonstrates the advantage of introducing the inductive bias to address low-correlations so that the models learn relations among a larger number of features better by filtering out less correlated features.
@@ 4.3 Utility (MLE)
The MLE results presented in Table 2 show that TabTreeFormer outperforms baseline models steadily regardless of the complexity of task, dataset, and ML model used. TTF-S achieves similar performance as the best-performing baselines (RTF and TabSyn), and TTF-M and TTF-L outperform them by a large margin of almost 40% in terms of relative errors in ML performance compared to real data. More MLE results can be found in Table 10 in Appendix D.2.
As an important foundation of tree-based models’ superior capability in capturing potentially low-correlations, rotational invariance of models is an aspect particularly worthwhile analyzing. As visualized in Fig. 9, rotationally invariant generators (e.g., CTAB-GAN+, CTGAN, GReaT, TVAE) are generally worse than non-rotationally invariant counterparts (e.g., REaLTabFormer and TabTreeFormer with a positional embedding in fixed order and TabSyn with VAE-encoding of data). Also, the advantage of TabTreeFormer over rotationally invariant baseline generators is more manifest on better-performing non-rotationally invariant downstream models (RF and XGB) than on rotationally invariant ones (LN). This again emphasizes the importance of applying non-rotationally invariant models to introduce inductive bias for low correlations.
@@ 4.4 Privacy (DCR)
Table 3: Privacy comparison of models. The first row shows the number of violations, and the second row shows the minimum $p$ value for privacy leakage risk. The closer the value to 0, the higher the risk of privacy leakage. Risky values are highlighted in red.
@@
As summarized in Table 3TTF-M and TTF-L are privacy-resilient in all the 10 datasets. TabTreeFormer-S has a privacy leakage in only one dataset, with the $p$-value very close to 0.05. In comparison, GReaT demonstrates severe privacy concerns in 5 out of 10 datasets. Although auto-regressive transformers suffer from privacy leakage concerns, TabTreeFormer and REaLTabFormer are generally privacy-preserving, validating the effectiveness of token masking in protecting privacy for tabular auto-regressive transformers. Raw $p$-values per experiment are presented in Table 11 in Appendix D.3.
4.5 Efficiency
The efficiency of TabTreeFormer is mainly demonstrated by its ability to achieve better performance with a smaller model size. TTF-M surpasses all baselines in most metrics (i.e., fidelity, MLE, and DCR) as shown in Section 4.2-4.4. Notably, it only has approximately 5M parameters while the most competitive counterparts have more than 20M (TabSyn) or even more than 80M parameters (GReaT and REaLTabFormer). Also, as shown in Table 4, there is no significant performance gap between TTF-M and TTF-L, although their sizes differ eight times. This demonstrates the capability of TabTreeFormer in maintaining high efficiency without significant compromise in effectiveness.
Table 5: Training time (s) comparison of TabTreeFormer and other auto-regressive transformer baselines on large datasets. The best values are highlighted in bold with an underscore, and the second best values are highlighted in bold without underscore.